OpenStudy (anonymous):

Choose the correct simplification of (8x + 5)(4x – 6).

OpenStudy (anonymous):

thank you for taking your time to answer this!

OpenStudy (unklerhaukus):

\[(8x + 5)(4x – 6)=32x^2-42x+20x-30\]

OpenStudy (unklerhaukus):

can you see what i have done?

OpenStudy (anonymous):

im afraid not these are the mutiple choices 32x2 + 28x + 30 32x2 + 28x − 30 32x2 − 28x + 30 32x2 − 28x − 30

OpenStudy (unklerhaukus):

ops eight times six is 48 not 42

OpenStudy (unklerhaukus):

\[(8x+5)(4x–6)=32x^2−48x+20x−30\] all you need to do now is simplify add the two x terms together

OpenStudy (anonymous):

32x2 + 28x − 30 so it is this <-----------

OpenStudy (unklerhaukus):

\[-48x+20x=-28x\]

OpenStudy (anonymous):

32x2 − 28x + 30 thats it right?

OpenStudy (unklerhaukus):

yes i hope you can see why

OpenStudy (anonymous):

yes thank you so much!